OpenStudy (anonymous):

can someone help me solve for x ... log of base 8 to the 16= x+1

OpenStudy (anonymous):

\[\log_{8}16 = x-1 \] use the "change-of-base" formula (log 16)/(log 8) = x-1 x = (log 16)/(log 8) + 1 x = 0.3333... or 1/3